1. Let X the number of tires on a randomly selected automobile that are un- derinflated. a). Which of the following three functions, p, f, and g, is a legitimate p.m.f for X, and why are the other two not allowed? p(x) 0.3 0.2 0.1 0.05 0.05 f(x) 0.4 0.1 0.1 0.1 0.3 g(x) 0.4 0.1 0.2 0.1 0.3
Answer #1

We know that sum of PMf over all possible values is equal to 1. So

Pr(x) = 1

In the first case, we have

〉 Pr(z) = (Pr(0)+PY ( 1 )+ Px (4)) = 0.3+0.2+0.1+0.05+0.05-0. T

So this is less than 1 and thus can't be a pmf.

Second, we have

〉 Pg(z) = (P, (0) Pr( 1)+ Pg(4)) = 0.4 + 0.1 + 0.1+0.1 + 0.3 = 1

So this is equal to 1 and is a pmf.

The third case is

〉 Pg(z) = (Pr(0)+Px (1 ) + Px(4)) = 0.4+0.1+0.2+0.1 + 0.3 = 1.1

This is greater than 1, thus it can't be a pmf.